# Asymptotically Optimal Lazy Lifelong Sampling-based Algorithm for Efficient Motion Planning in Dynamic Environments

### v2

@@ Abstract
The paper introduces an asymptotically optimal lifelong sampling-based path planning algorithm that combines the merits of lifelong planning algorithms and lazy search algorithms for rapid replanning in dynamic environments where edge evaluation is expensive. By evaluating only sub-path candidates for the optimal solution, the algorithm saves considerable evaluation time and thereby reduces the overall planning cost. It employs a novel informed rewiring cascade to efficiently repair the search tree when the underlying search graph changes after evaluation. Simulation results demonstrate that the algorithm outperforms various state-of-the-art sampling-based planners in addressing both static and dynamic motion planning problems.
@@
Motion planning, lifelong planning, dynamic environments, sampling-based, asymptotically optimal, lazy search
@@ I Introduction
Real-time motion planning is crucial for robots to navigate safely in partially known or completely unknown environments. In such scenarios, robots must interleave planning with action executions to ensure that their plans remain viable and safe, accounting for dynamic obstacles. Efficient planning algorithms are necessary in these cases. These algorithms should be capable of updating previous solutions based on new environment information in a short time.
@@
Early researches proposed that reusing previous search can significantly enhance planning efficiency [1, 2]. The strategy, commonly known as lifelong planning in literature, was combined with graph-based path planning techniques, such as A [3]. In particular, Lifelong Planning A (LPA) [4] and its variants D/D Lite [5], repair the same search graph that approximates the problem domain throughout the entire navigation process. Any inconsistency resulting from edge changes in the graph is identified, and the inconsistencies are efficiently propagated to the relevant parts of the search graph to repair an A-like solution. It ensures efficient adaptation to the dynamic nature of the problem domain. However, LPA is only resolution optimal and complete with respect to the particular underlying graph used. In other words, they almost surely find sub-optimal paths with respect to the robot and environment. Improving completeness/optimality requires recomputing the search graph, which is infeasible in applications.
@@
Sampling-based planning methods, such as RRT [6], RRT [7], Informed RRT [8], in contrast, avoid an explicit offline construction of the problem domain and refines an asymptotically dense Random Rapidly-explored Graph (RRG) directly in the robot’s state-space. These methods aim to find asymptotically optimal paths and probabilistically complete in terms of the robot and its environment. However, a significant drawback of these methods is their inability to efficiently handle graph changes, as they require replanning from scratch whenever a change is detected. To address this limitation, previous works have explored reusing previous search trees to improve the replanning efficiency of sampling-based planning methods [9, 10, 11, 12, 13]. Unfortunately, these methods do not guarantee the quality of their solutions. In contrast, asymptotically optimal lifelong sampling-based methods [14, 15] stand out from previous approaches by offering asymptotically optimal solutions. They use a rewiring cascade to remodel the search tree around dynamic obstacles promptly. These methods offer significant benefits, particularly in applications that require high solution quality.
@@
Unfortunately, the existing asymptotically optimal lifelong sampling-based methods primarily focused on enhancing solution quality but need to be more indifferent to the number of edge evaluations. Specifically, they must iteratively evaluate all edges to check their feasibility before rewiring, irrespective of their relevance to the solution. In motion planning problems, edge evaluation typically consists of computationally-cost procedures, such as solving two-point boundary value problems and conducting dynamic collision checks. As a result, the cost of edge evaluation dominates the overall replanning process, especially for motion planning problems in high-dimensional configuration spaces, where the RRG must be dense enough to contain a solution.
@@
The lazy search technique has been introduced to mitigate the problem of excessive edge evaluations [16]. It assumes that a heuristic function can efficiently compute a lower bound on edge weight and uses the lower bound to estimate whether an edge has the potential to improve the solution. This approach defers the evaluation of the actual cost of an edge until it is deemed pertinent to the solution. Various sampling-based algorithms [17, 18, 19] have exploited the lazy search strategy to enhance their planning efficiency. Regrettably, none can achieve both the replanning efficiency of lifelong planning techniques and the edge evaluation efficiency of lazy search techniques.
@@
In this paper, we propose an asymptotically optimal lifelong sampling-based algorithm, Lazy Lifelong Planning Tree (LLPT), for online motion planning in dynamic environments where edge evaluation is expensive. The algorithm integrates the lazy search framework and the lifelong planning framework. Specifically, LLPT approximates the problem domain by an RRG with non-overestimating heuristic edges. The actual edge evaluations of the RRG are restricted to those with the potential to be part of the solution. Whenever the RRG changes (e.g., the actual cost of some edges are evaluated or the RRG is refined), a novel informed rewiring cascade is performed to quickly repair the search tree that spans the RRG to reflect the new information and update the solution. We show that LLPT outperforms several state-of-the-art planning algorithms (RRT[6], Informed RRT[8], BIT[19], RT-RRT[14], RRTX[15]) over a set of motion planning problems, including both static and dynamic ones.
@@ II-A Lifelong Sampling-based Algorithms
Over the past decades, lifelong sampling-based planning for dynamic motion planning problems has evolved through various iterations. Rather than discarding all previous samples like brute-force replanning, some algorithms selectively eliminate invalid search tree branches obstructed by dynamic obstacles [9, 10]. They then incrementally densify the RRG until a new solution is derived. More efficient replanning schemes only remove invalid edges of the spanning tree and preserve the separate branches for further planning. For instance, LRF retains a forest of spanning trees that are dynamically split, regrown, and merged to adapt to obstacles and robot movements [11]. Similarly, MP-RRT retains obstructed branches and actively seeks to reconnect them to the root tree, bridging the gap between the start and goal configurations [12]. RRTFND reduces replanning effort by only reconnecting or regrowing functional disconnected branches of the previous solution path [13].
However, the aforementioned approaches are lack of formal guarantees on solution optimality. Distinguished from previous works, RT-RRT[14] and RRTX[15] were designed to address dynamic shortest-path planning problems (DSPP), offering an asymptotically optimal solution given sufficient replanning time. These methods leverage rewiring cascades to adapt the search tree and navigate around dynamic obstacles. In particular, RT-RRT rewires the entire search tree structure starting from the root, which can be computationally expensive when there are many nodes in the tree. On the other hand, RRTX uses a local rewiring cascade, which focuses only on rewiring the parts of the search tree affected by dynamic changes. This makes it more efficient in scenarios where the environment experiences minor alterations.
@@
II-B Sampling-Based Planning with Lazy Collision Detection
The lazy sampling-based planning is based on the assumption that a heuristic function can estimate a lower bound on edge costs effectively. The lower bound is used to avoid unnecessary edge evalution. Lazy PRM and Lazy RRG [17] build a roadmap of feasible configurations like PRM and RRG [6], but with the difference that edges are not immediately checked for collision. Only edges that are on a better candidate path to the goal are checked for collision. If an edge is found to collide, it is removed from the roadmap, and the candidate path is updated. Similarly, Lower Bound Tree-RRT (LBT-RRT) [18] proposed the lower bound graph $\mathcal{G}_{lb}$, an RRG with lazily-evaluated edges, to approximate the problem domain. $\mathcal{G}_{lb}$ poses a lower bound invariant to the nodes in the search tree, i.e., the solution cost of the nodes in the search tree is always smaller than their solution cost via $\mathcal{G}_{lb}$. An edge is only evaluated if its lazy insertion into $\mathcal{G}_{lb}$ causes the violation against the lower bound invariant. Batch Informed Tree (BIT) [19] inserts a batch of samples concurrently into the search graph with lazily evaluated edges and prioritizes the evaluation of the new edges with minor lazily-evaluated solution cost. This mechanism boosts evaluation efficiency, albeit at the expense of queuing up edges in the process.
Our work brings together lifelong sampling-based algorithms and lazy search techniques to harvest replanning and edge evaluation efficiency when solving DSPP. Our algorithm outperforms existing lifelong and lazy sampling-based algorithms in terms of planning efficiency for both static and dynamic planning problems.
@@ III Lazy Lifelong Planning Tree
In this section, we first introduce the notations that will be used throughout the paper. In Section III-B, we highlight the informed rewiring cascade, a crucial component of LLPT that efficiently rewires the underlying graph upon detecting modifications. We then present the overall algorithm framework, with a particular focus on the lazy search strategy that is used to delay collision check. Finally, we describe the graph densification strategy used to enhance the solution.
@@ III-A Notations and Problem Definitions
Assume that a robot begins its journey from $v_{start}$ to $v_{goal}$. An optimal path planning problem is to find a path $\pi^{*}(v_{start},v_{goal})$, an ordered set of distinct vertices that connects $v_{start}$ and $v_{goal}$ and minimizes the total cost to traverse this path. LLPT solves this problem by constructing a search graph $\mathcal{G}=(V,E)$ with lazily-evaluated edges within the robot’s configuration space $\mathcal{X}$, where $V$ represents the node set and $E\subseteq V\times V$ denotes the edge set. To clarify, each node $v\in V$ implies the corresponding configuration $x\in\mathcal{X}$ and can be directly input into distance functions. For each node, LLPT stores its neighbor nodes in the memory for future reference. The search tree of $\mathcal{G}$ is denoted by $\mathcal{T}=(V_{\mathcal{T}},E_{\mathcal{T}})$, rooted at $v_{goal}$, where $V_{\mathcal{T}}\subseteq V$ and $E_{\mathcal{T}}\subseteq E$. As mentioned in [15], Rooting $\mathcal{T}$ in $v_{goal}$ allows the same tree and root to be used for the entire navigation, avoiding rewiring the tree from the root while the robot moves.
@@
The solution path should minimize the total weight, i.e., $\pi^{*}(v_{start},v_{goal})=\arg\min_{\pi\in\mathcal{G}}\sum_{e\in\Pi}w(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
@@ 8             RemoveFromTree($v$);
9             $lmc(v)\leftarrow\infty$;
10             $E_{collided}\xleftarrow{+}\{e\}$, $V_{collided}\xleftarrow{+}\{v\}$;
11
12      $E_{eval}\xleftarrow{+}\{e\}$;
13
14return $V_{collided}$;
Algorithm 2 EvaluateEdge($\pi$, $\alpha$)
1 $lmc(v)\leftarrow\infty$, $V_{\mathcal{T}}\xleftarrow{-}v$, $\mathcal{Q}$.update($v$);
2 for $u\in\textnormal{{Child(}}\textnormal{\emph{$v$}}\textnormal{{)}}$ do

### v3

@@ Abstract
The paper introduces an asymptotically optimal lifelong sampling-based path planning algorithm that combines the merits of lifelong planning algorithms and lazy search algorithms for rapid replanning in dynamic environments where edge evaluation is expensive. By evaluating only sub-path candidates for the optimal solution, the algorithm saves considerable evaluation time and thereby reduces the overall planning cost. It employs a novel informed rewiring cascade to efficiently repair the search tree when the underlying search graph changes. Theoretical analysis indicates that the proposed algorithm converges to the optimal solution as long as sufficient planning time is given. Planning results on robotic systems with $\mathbb{SE}(3)$ and $\mathbb{R}^{7}$ state spaces in challenging environments highlight the superior performance of the proposed algorithm over various state-of-the-art sampling-based planners in both static and dynamic motion planning tasks. The experiment of planning for a Turtlebot 4 operating in a dynamic environment with several moving pedestrians further verifies the feasibility and advantages of the proposed algorithm.
Note to Practitioners
Fast and adaptive replanning is essential for robotic systems operating in dynamic environments with limited information and computational resources. While sampling-based planners are well-suited to such settings due to their ability to incrementally construct paths, their performance is often hindered by costly edge collision checking. In this work, we propose a lifelong sampling-based planning algorithm that maintains and reuses a global roadmap over time, integrating a lazy edge evaluation strategy that defers collision checks until necessary. This approach enables efficient, continual replanning without rebuilding the search structure from scratch. Simulations and experiments in dynamic scenarios demonstrate improved planning speed and responsiveness, making the method suitable for real-time robotic applications.
@@
Motion planning, lifelong planning, dynamic environments, asymptotically optimal, lazy search
@@ I Introduction
Autonomous robots frequently operate in environments that are partially or entirely unknown. For instance, a service robot navigating a crowded shopping mall or pedestrian area, a mobile robot using sensors to move outdoors without a complete roadmap, or multiple warehouse robots collaborating without predefined trajectories. In such scenarios, robots must quickly update their solutions based on new sensory data to ensure robust and responsive autonomy.
@@
Early research proposed that reusing the previous searches can significantly enhance planning efficiency [1, 2]. The strategy, commonly known as lifelong planning in literature, was combined with graph-based path planning techniques, such as A [3]. In particular, Lifelong Planning A (LPA) [4] and its variants D/D Lite [5] repair the same search graph that approximates the problem domain throughout the entire navigation process. Any inconsistency resulting from edge changes in the graph is identified, and the inconsistencies are efficiently propagated to the relevant parts of the search graph to repair an A-like solution. It ensures efficient adaptation to the dynamic nature of the problem domain. However, LPA is only resolution-optimal and -complete with respect to the particular underlying graph used. In other words, they almost surely find sub-optimal paths with respect to the robot and environment. Improving completeness/optimality requires recomputing the search graph, which is infeasible in applications.
Figure 1: Graphical illustrations of LLPT finding the shortest path from start to goal in the presence of a dynamic obstacle. The gray dashed lines represent unevaluated but potentially valid edges, while the red dashed lines denote unevaluated and potentially invalid edges. The black lines indicate evaluated and valid edges, and the solid red lines denote evaluated and invalid (i.e., colliding) edges. (a) The configuration space is initially spanned by a search tree composed of unevaluated edges. (b) A solution path from start to goal is found and verified to be collision-free. (c) A dynamic obstacle moves, potentially altering the collision status of nearby edges. (d) The edges along the previously found solution path are re-evaluated, and a collision is detected on one of them. (e) The search tree is incrementally rewired, and a new solution path is identified. (f) The new path is evaluated, and another colliding edge is discovered. (g) The tree is rewired again, and the planner falls back to a suboptimal solution with higher cost. (h) All edges along the new solution are finally evaluated and found to be collision-free. The solution is updated accordingly.
@@
Asymptotically-optimal sampling-based planning methods, such as Rapidly-exploring Random Tree (RRT) [6], RRT [7], Informed RRT [8], in contrast, avoid an explicit offline construction of the problem domain and refine an asymptotically dense space-filling tree directly in the robot’s state-space. These methods aim to find asymptotically optimal paths in terms of the robot and its environment. However, a significant drawback of these methods is their inability to efficiently handle environment changes, as they require replanning from scratch whenever a change is detected. To address this limitation, previous works have explored reusing previous search trees to improve the replanning efficiency of sampling-based planning methods [30, 31, 32, 33, 34]. Unfortunately, these methods do not guarantee the quality of their solutions. In contrast, asymptotically optimal lifelong sampling-based methods [35, 36, 37] stand out from previous approaches by offering asymptotically optimal solutions. They use a rewiring cascade to promptly remodel the search tree around dynamic obstacles. These methods offer significant benefits, particularly in applications that require high solution quality.
@@
Unfortunately, the existing asymptotically optimal lifelong sampling-based methods primarily focused on enhancing solution quality but need to be more indifferent to the number of edge evaluations. Specifically, they must iteratively evaluate all edges of the underlying search graph to check their feasibility before rewiring, irrespective of their relevance to the solution. In motion planning problems, edge evaluation typically consists of computationally-intensive procedures, such as solving two-point boundary value problems and conducting dynamic collision checks. As a result, the cost of edge evaluation dominates the overall replanning process, especially for motion planning problems in high-dimensional configuration spaces, where the underlying search graph must be dense enough to contain a solution.
@@
The lazy search technique has been introduced in [46] to mitigate the problem of excessive edge evaluations. It assumes that a heuristic function can efficiently compute a lower bound on edge weight and uses the lower bound to estimate whether an edge has the potential to improve the solution. This approach defers the evaluation of the actual cost of an edge until it is deemed pertinent to the solution. Various sampling-based algorithms [47, 48, 49] have exploited the lazy search strategy to enhance their planning efficiency. Regrettably, none can achieve both the replanning efficiency of lifelong planning techniques and the edge evaluation efficiency of lazy search techniques.
@@
In this paper, we propose an asymptotically optimal lifelong sampling-based algorithm, Lazy Lifelong Planning Tree (LLPT), for online motion planning in environments where edge evaluation is expensive. The algorithm integrates the lazy search and the lifelong planning frameworks. Specifically, LLPT approximates the problem domain by a Rapidly-exploring Random Graph (RRG) with non-overestimating heuristic edges. The RRG is reused and refined throughout the entire planning procedure until the robot reaches the goal. LLPT searches for solutions by spanning the RRG with a Shortest-Path Tree (SPT). The actual edge evaluations of the SPT are restricted to those with the potential to be part of the solution. Whenever changes occur (e.g., the actual cost of some edges is evaluated or the RRG is refined), a novel informed rewiring cascade is performed to quickly repair the SPT to reflect the new information and update the solution.
This paper extends our previous work published at IROS 2024 [51]. We provide a comprehensive theoretical analysis, including an expanded proof of the probabilistic completeness and asymptotic optimality of LLPT (see Section IV). We also compare LLPT with a broader range of sampling-based motion planning algorithms on various high-dimensional planning problems, including both static and dynamic scenarios, to further demonstrate its advantages (see Section V). The same section presents real-world experiments conducted on a Turtlebot 4 platform, validating the feasibility and efficiency of the proposed algorithm. Furthermore, we evaluate the performance of LLPT in highly dynamic environments and discuss potential directions for future improvement in the discussion part.
@@ II-A Lifelong Sampling-based Algorithms
The concept of reusing previous search efforts is widely adopted by various planners to enhance their planning efficiency. Anytime motion planners [9, 8, 10, 6, 7] refine their solutions by adding new samples to existing search trees, leveraging previously found solution paths. However, these methods only reuse search efforts within a single query and must restart when the robot’s operating environment or its start and goal configurations change. Multi-query sampling-based planners, such as Probability Roadmap (PRM) [11] and RRG [6], address this limitation by efficiently solving multiple start-goal queries in static environments. They achieve this by reusing previously acquired knowledge of valid edges, which allows for more efficient solutions to subsequent queries.
Some sampling-based motion planners focus on reusing planning results to manage environmental variations effectively. For instance, certain approaches build trajectory libraries [12] that store previous successful plans, guiding future searches based on handcrafted or learned heuristics, such as roadmap similarity [13], collision probability models [14], or the proximity of start-goal configurations [15]. These trajectory libraries have proven effective for controlling underactuated and high-dimensional systems [16, 17]. However, this approach can suffer from limited dataset sizes or excessive memory usage. The Thunder framework [18] enhances trajectory library-based planners by storing experiences in a sparse graph rather than in individual paths. It reduces redundancy and increases opportunities for path reuse.
Additionally, significant efforts have been made to learn latent representations of configuration spaces from past experiences. This informs future queries by biasing sample sequences toward regions likely to contain a solution path [19, 20, 21]. Some approaches leverage neural networks to improve the accuracy of these latent representations [56, 57, 58]. Others view biased sampling as a decision-making process, employing reinforcement learning methods to learn optimal sampling policies online [27, 28, 29]. However, these learning-based approaches often require extensive offline or online training to ensure the effectiveness of the sampler parameters.
Different from previous approaches, some accelerate replanning by rewiring the out-of-date search trees to fully exploit previous search efforts. DRRT and GRIP [30, 31] selectively eliminate invalid search tree branches obstructed by dynamic obstacles rather than discarding all previous samples as brute-force replanning. They then grow the search tree until a new solution is derived. More efficient replanning schemes only remove invalid edges of the spanning tree and preserve the separate branches for further planning. For instance, LRF retains a forest of spanning trees that are dynamically split, regrown, and merged to adapt to obstacles and robot movements [32]. Similarly, MP-RRT retains obstructed branches and actively seeks to reconnect them to the root tree, bridging the gap between the start and goal configurations [33]. RRTFND reduces replanning effort by only reconnecting or regrowing functional disconnected branches of the previous solution path [34].
However, the aforementioned approaches are lack of formal guarantees on solution optimality. Distinguished from previous works, RT-RRT[35], RRTX[36], and FAT[37] were designed to address dynamic shortest-path planning problems (DSPP), offering an asymptotically optimal solution given sufficient replanning time. These methods leverage rewiring cascades to adapt the search tree and navigate around dynamic obstacles. In particular, RT-RRT rewires the entire search tree structure starting from the root, which can be computationally expensive when there are many nodes in the tree. On the other hand, RRTX uses a local rewiring cascade, which focuses only on rewiring the parts of the search tree affected by dynamic changes, rendering it more efficient in scenarios where the environment experiences minor alterations. FAT further optimizes the rewiring process of RRTX by constraining expansions to promising samples that are likely to improve solution costs.
@@
II-B Path Planning with Lazy Collision Detection
The issue of expensive edge evaluations is particularly significant in robotics. While numerous studies aim to reduce collision checks by biasing sampling toward free regions [38, 39], it is common for the motion between pairs of free samples to be infeasible, especially in environments with narrow passages or dense obstacles. One effective strategy is to select paths that minimize the likelihood of collisions. Pan and Manocha develop incremental collision probability models by which only the paths that are estimated to be feasible are evaluated [40]. This concept has also been explored within a Bayesian reinforcement learning framework [41]. Choudhury et al. [42] frame edge selection as a decision region determination problem, choosing edges that reduce uncertainty about the robot’s operating environment. STROLL [43] formulates the edge selection problem as a Markov Decision Process (MDP), learning an edge evaluation policy by imitating the strategy of LazySP [44]. Additionally, Sung and Stone [45] prioritize the evaluation of edges located on the cut of the underlying graph, as these edges are more closely related to potential solutions.
On the other hand, some works assume that a heuristic function can estimate a lower bound on edge costs effectively and guide collision checks by estimated solution length. Lazy PRM and Lazy RRG [47] build a roadmap of feasible configurations like PRM and RRG [6], but with the difference that edges are not immediately checked for collision. Only edges that are estimated by the heuristic function to belong to a better candidate solution path are checked for collision. If an edge is found to collide, it is removed from the roadmap, and the candidate path is updated. Similarly, Lower Bound Tree-RRT (LBT-RRT) [48] proposed the lower bound graph $\mathcal{G}_{lb}$, an RRG with lazily-evaluated edges. $\mathcal{G}_{lb}$ poses a lower bound invariant to the nodes in the search tree, i.e., the solution cost of the nodes in the search tree is always smaller than their solution cost via $\mathcal{G}_{lb}$. An edge is only evaluated if its lazy insertion into $\mathcal{G}_{lb}$ causes the violation against the lower bound invariant. Batch Informed Tree (BIT) [49] inserts a batch of samples concurrently into the search graph with lazily evaluated edges and prioritizes the evaluation of the new edges with minor lazily-evaluated solution cost.
Our work brings together lifelong sampling-based algorithms and lazy search techniques to simultaneously harvest replanning and edge evaluation efficiency when solving DSPP. Our algorithm outperforms existing lifelong and lazy sampling-based algorithms in terms of planning efficiency for both static and dynamic planning problems.
@@ III Lazy Lifelong Planning Tree
Figure 2: System overview of the proposed replanning module integrated with a complete robot navigation system.
In this section, we describe the proposed lifelong planning framework. The system overview is presented in 2. In the following subsection, we first introduce the notations that will be used throughout the paper. In Section III-B, we highlight the informed rewiring cascade, a crucial component of LLPT that efficiently rewires the underlying graph upon detecting graph modifications. We then present the overall algorithm framework, with a particular focus on the lazy search strategy that is used to delay collision checks. Finally, we describe the graph densification strategy used to enhance the solution.
@@ III-A Notations and Problem Definitions
Assume that a robot begins its journey from $v_{start}$ to $v_{goal}$. An optimal path planning problem is to find a path $\pi^{*}(v_{start},v_{goal})$, an ordered set of distinct vertices that connects $v_{start}$ and $v_{goal}$ and minimizes the total cost to traverse this path. LLPT solves this problem by constructing a search graph $\mathcal{G}=(V,E)$ with lazily-evaluated edges within the robot’s configuration space $\mathcal{X}$, where $V$ represents the node set and $E\subseteq V\times V$ denotes the edge set. To clarify, each node $v\in V$ implies the corresponding configuration $x\in\mathcal{X}$ and can be directly input into distance functions. For each node, LLPT stores its neighbor nodes in memory for future reference. The search tree of $\mathcal{G}$ is denoted by $\mathcal{T}=(V_{\mathcal{T}},E_{\mathcal{T}})$, rooted at $v_{goal}$, where $V_{\mathcal{T}}\subseteq V$ and $E_{\mathcal{T}}\subseteq E$. As mentioned in [36], rooting $\mathcal{T}$ in $v_{goal}$ allows the same tree and root to be used for the entire navigation, avoiding rewiring the tree from the root while the robot moves.
@@
The goal of our planner is to find the optimal solution path $\pi^{*}$ that minimizes the total true edge weight, i.e., $\pi^{*}(v_{start},v_{goal})=\arg\min_{\pi\in\mathcal{G}}\sum_{e\in\Pi}w(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
@@ 4       $E_{uneval}\leftarrow$ first $\alpha$ edges in $E_{uneval}$ closest to $v_{goal}$;
8             $E_{collided}\xleftarrow{+}\{e\}$, $V_{collided}\xleftarrow{+}\{v\}$;
9
10      $E_{eval}\xleftarrow{+}\{e\}$;
11
12return $V_{collided}$;
Algorithm 2 EvaluateEdge($\pi$, $\alpha$)
@@ 1 RemoveFromTree($v$);
2 $lmc(v)\leftarrow\infty$, $\mathcal{Q}$.update($v$);
3 for $u\in$ Child($v$) do